Give the equation of any vertical, horizontal, or oblique asymptote for the graphs of each ration function in #1.
f(x)=(x^2-1)/(x+3)

Answer by lwsshak3(11628)

Give the equation of any vertical, horizontal, or oblique asymptote for the graphs of each ration function in #1.
f(x)=(x^2-1)/(x+3)
**
Since degree of denominator is one less than degree of numerator, there is an oblique asymptote.
By long division, divide numerator by denominator. The equation of the oblique asymptote is the quotient, ignoring the remainder=x-3
equation of the oblique asymptote: y=x-3
..
To find vertical asymptote, set denominator=0, then solve for the x-values which make the function undefined.
x+3=0
x≠-3
equation of vertical asymptote:x=-3
..
horizontal asymptotes: none